The amount of time a service technician needs to change the oil in a car is uniformly distributed between 10 and 23 minutes. Let X = the time needed to change the oil on a car. Write the distribution.

Find P (x > 20).

Find the 65th percentile.
Answer #1

Solution :

Given that,

a = 10

b = 23

a) P(x < c) = (c - a) / (b - a)

P(x < 20) = (20 - 10) / (23 - 10)

P(x < 20) = 0.7692

b) 65 th PERCENTILE

= a + (b - a)p

= 10 + (23 - 10)0.65

= 18.45 minutes